# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ 4.3 Marginal constraints in risk aggregation
Using $\boldsymbol{\gamma}\prec\boldsymbol{w}$, there exists $\Lambda\in\mathcal{Q}_{n}$ such that $(F_{1}^{-1},\dots,F_{n}^{-1})=\Lambda(G_{1}^{-1},\dots,G_{n}^{-1})$, denoted by $\mathbf{F}=\Lambda\otimes\mathbf{G}$. By Theorem 3 of Chen et al. (2022), for $\alpha\in(0,1)$, we have
$\sup_{G\in\mathcal{D}_{n}(\mathbf{G})}q_{\alpha}^{-}(G)\leqslant\sup_{G\in\mathcal{D}_{n}(\Lambda\otimes\mathbf{G})}q_{\alpha}^{-}(G)=\sup_{G\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(G)$
for $\alpha\in(0,1)$. Hence, it follows from (i) of Corollary 1 that $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})\leqslant q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in(0,1)$. Using the fact that both $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $F_{\mathcal{D}_{n}(\mathbf{G})}^{-}$ are right-continuous for $\alpha\in(0,1)$, we have $F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)\geqslant F_{\mathcal{D}_{n}(\mathbf{G})}^{-}(x)$ for all $x\in\mathbb{R}$. By the monotonicity of $\rho$, we have $\rho(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})\leqslant\rho(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Moreover, by Theorem 4, we have $\overline{R}_{\rho}(\boldsymbol{\gamma})=\rho(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ and $\overline{R}_{\rho}(\boldsymbol{w})=\rho(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Hence, we have $\overline{R}_{\rho}(\boldsymbol{w})\leqslant\overline{R}_{\rho}(\boldsymbol{\gamma}).$
Note that Proposition 11 shows that for the assets with same marginal distributions $F\in\mathcal{M}_{D}\cup\mathcal{M}_{I}$, more diversified portfolio has higher risk under dependence uncertainty if $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$ is applied. This is not surprising as some similar conclusion is shown in Proposition 8 of Chen et al. (2022) for quantiles. Note that if $\Lambda$ is a constant, $\tilde{q}_{\Lambda}^{-}$ boils down to quantiles. Therefore, the conclusion in Proposition 11 can be viewed as an extension of Proposition 8 of Chen et al. (2022). Moreover, Proposition 11 also implies that under those assumptions, the optimal portfolio positions are the ones with only single asset.
@@ 5 Conclusion
This paper summaries some properties of $\Lambda$-quantiles and explores distributionally robust models of $\Lambda$-quantiles. We obtain robust $\Lambda$-quantiles on general uncertainty sets, showing that obtaining the robust $\Lambda$-quantiles relies on finding the extreme probabilities over the same uncertainty sets. This finding is surprising and significantly simplifies the problem, enabling us to utilize many existing results on the robust quantiles from the literature. We provide closed-form solutions for the uncertainty sets characterized by three different constraints: (i) moment constraints; (ii) probability constraints via Wasserstein distance; and (iii) marginal constraints in risk aggregation. Those results are applied to optimal portfolio selection under model uncertainty, and can also be extended to other problems such as optimal reinsurance, which will be discussed in future.
@@ Acknowledgments
The research of Xia Han is supported by the National Natural Science Foundation of China (Grant No. 12301604, 12371471) and the Fundamental Research Funds for the Central Universities, Nankai University (Grant No. 63231138).

### v2

@@ 4.3 Marginal constraints in risk aggregation
Using $\boldsymbol{\gamma}\prec\boldsymbol{w}$, there exists ${A}\in\mathcal{Q}_{n}$ such that $(F_{1}^{-1},\dots,F_{n}^{-1})={A}(G_{1}^{-1},\dots,G_{n}^{-1})$, denoted by $\mathbf{F}=A\otimes\mathbf{G}$. By a slight extension of Theorem 3 of Chen et al. (2022), for $\alpha\in[t,1)$, we have
$\sup_{G\in\mathcal{D}_{n}(\mathbf{G})}q_{\alpha}^{-}(G)\leqslant\sup_{G\in\mathcal{D}_{n}({A}\otimes\mathbf{G})}q_{\alpha}^{-}(G)=\sup_{G\in\mathcal{D}_{n}(\mathbf{F})}q_{\alpha}^{-}(G).$
Hence, it follows from (i) of Corollary 1 that $q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{G})}^{-})\leqslant q_{\alpha}^{-}(F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ for $\alpha\in[t,1)$. By Proposition 11, we have $t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-}$ and $t\vee F_{\mathcal{D}_{n}(\mathbf{G})}^{-}$ are continuous over $\mathbb{R}$. Hence, we have $t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-}(x)\leqslant t\vee F_{\mathcal{D}_{n}(\mathbf{G})}^{-}(x)$ for all $x\in\mathbb{R}$. By the monotonicity of $\rho$, we have $\rho(t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-})\geqslant\rho(t\vee F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Moreover, by Theorem 4, we have $\overline{R}_{\rho}(\boldsymbol{\gamma})=\rho(t\vee F_{\mathcal{D}_{n}(\mathbf{F})}^{-})$ and $\overline{R}_{\rho}(\boldsymbol{w})=\rho(t\vee F_{\mathcal{D}_{n}(\mathbf{G})}^{-})$. Hence, we have $\overline{R}_{\rho}(\boldsymbol{w})\leqslant\overline{R}_{\rho}(\boldsymbol{\gamma}).$
Note that Proposition 12 shows that for the assets with same marginal distributions $F\in\mathcal{M}_{t,D}^{+}\cup\mathcal{M}_{t,I}^{+}$ with $0\leqslant t<\alpha_{\Lambda}^{-}$, more diversified portfolio has higher risk under dependence uncertainty if $\rho=\tilde{q}_{\Lambda}^{+}$ with decreasing $\Lambda$ or $\rho=\tilde{q}_{\Lambda}^{-}$ is applied. This is not surprising as some similar conclusions are shown in Proposition 8 of Chen et al. (2022) for quantiles under the assumption that the marginal distribution $F$ has a monotone density. Note that if $\Lambda$ is a constant, $\tilde{q}_{\Lambda}^{-}$ boils down to quantiles, and if $t=0$, then $F$ has a monotone density. Therefore, the conclusion in Proposition 12 can be viewed as an extension of Proposition 8 of Chen et al. (2022). Moreover, Proposition 12 also implies that under those assumptions, the optimal portfolio positions are the ones with only a single asset.
@@ 5 Conclusion
This paper summarizes some properties of $\Lambda$-quantiles and explores distributionally robust models of $\Lambda$-quantiles. We obtain robust $\Lambda$-quantiles on general uncertainty sets under some conditions, showing that obtaining the robust $\Lambda$-quantiles relies on finding the extremal distributions over the same uncertainty sets. This finding significantly simplifies the problem, enabling us to utilize many existing results on the robust quantiles from the literature. We provide closed-form solutions for the uncertainty sets characterized by three different constraints: (i) moment constraints; (ii) probability constraints via the Wasserstein distance; and (iii) marginal constraints in risk aggregation. These results are applied to optimal portfolio selection under model uncertainty, and can also be extended to other problems such as optimal reinsurance, which will be discussed in the future.
@@ Acknowledgments
The authors thank the editor, an associate editor, and two anonymous referees for their helpful comments. The research of Xia Han is supported by the National Natural Science Foundation of China (Grant Nos. 12301604, 12371471, and 12471449). Peng Liu is the corresponding author.
Data availability statement
N/A. This is a theoretical paper and it does not make use of real data.
Conflict of interest disclosure
None of the authors have a conflict of interest to disclose.